TEX document #4

\overline{x} = x_{cm} = \frac{∫∫∫_{E} ρ(x, y, z) x dV}{M} ,

\overline{y} = \frac{∫∫∫_{E} ρ(x, y, z) y dV}{M} , \overline{z} = \frac{∫∫∫_{E} ρ(x, y, z) z dV}{M} .

\overline{x} = \frac{∫∫∫_{E} x dV}{∫∫∫_{E} dV} = \frac{∫∫∫_{E} x dV}{Volume of E} , etc.